# local non-bossiness

### v4

@@ A $\mu$-improving cycle is a tuple of different students $(i_{1},\ldots,i_{r})$ such that
In particular, this property implies that the converse of local non-bossiness is satisfied in this context: if the colleagues of some student do not change when her preferences are modified, then her school does not change too. $\Box$
Notice that, ${\rm{DA}}$ is not the only locally non-bossy and stable mechanism that can be defined in ${\mathcal{P}}$. Indeed, the school-optimal stable mechanism also satisfies these properties, because it is non-bossy (Afacan and Dur, 2017). Furthermore, the arguments made in the proof of Theorem 1 can be easily adapted to provide a direct proof that the school-optimal stable mechanism is locally non-bossy. Indeed, since this mechanism assigns each student to the worst school she can reach in a stable matching, it is sufficient to repeat the proof focusing on $\mu$-worsening cycles (i.e., tuples $(i_{1},\ldots,i_{r})$ such that $(i_{r},\ldots,i_{1})$ forms a $\mu$-improving cycle).
It is well-known that ${\rm{DA}}$ is the only mechanism in ${\mathcal{P}}$ that is both stable and strategy-proof (Alcalde and Barberà, 1994). Since locally group strategy-proofness is stronger than strategy-proofness, and locally non-bossiness and strategy-proofness ensure locally group strategy-proofness (Lemma 1), the following result is a direct consequence of Theorem 1.
@@ 4. A characterization of DA via local non-bossiness
In this section, we provide a characterization of DA using axioms that do not involve priorities. Our result extends the characterization of DA given by Ehlers and Klaus (2016, Theorem 1) from one-to-one matching problems to many-to-one contexts.
To maintain the generality of Ehlers and Klaus’ (2016) context, we analyze a variable-population school choice model. Hence, given a population $\overline{N}$ of students and a set $S$ of schools with capacities $q=(q_{s})_{s\in S}$, we consider mechanisms that depend only on the set $N\subseteq\overline{N}$ of students in the market and their preferences $(P_{i})_{i\in N}$.
To simplify notations, we consider ${\mathcal{P}}={\mathcal{L}}^{\overline{N}}$ as the preference domain, although only the preferences of the agents in the market will be relevant at the moment of determine a matching. Let ${\mathcal{N}}$ be the collection of non-empty subsets of $\overline{N}$. Given $N\in{\mathcal{N}}$, let $\mathcal{M}(N)$ be the set of matchings compatible with $[N,S,q]$. That is, $\mathcal{M}(N)$ is the family of functions $\mu:N\rightarrow S\cup\{s_{0}\}$ such that |$|\{i\in N:\mu(i)=s\}|\leq q_{s}$ for all $s\in S$. Let $\mathcal{M}^{*}=\bigcup_{N\in{\mathcal{N}}}\mathcal{M}(N)$. For each $i\in\overline{N}$ and $P_{i}\in{\mathcal{L}}$, denote by $A(P_{i})=\{s\in S:sP_{i}s_{0}\}$ the set of schools that are admissible under $P_{i}$.
A mechanism is a function $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ that associates to each $(N,P)\in{\mathcal{N}}\times\mathcal{P}$ a matching in ${\mathcal{M}}(N)$ that only depends on $[N,S,q,(P_{i})_{i\in N}]$.
@@
A mechanism $\Phi$ is weakly WrARP when for all $N,N^{\prime}\in{\mathcal{N}}$, $P\in{\mathcal{P}}$, and $s\in S$ such that $|N|=|N^{\prime}|=q_{s}+1$ and $A(P_{k})=\{s\}$ for all $k\in N\cup N^{\prime}$,
$\quad\Big{[}\,i,j\in N\cap N^{\prime},\,\,\quad i\in\Phi_{s}(N,P),\,\,\quad j\in\Phi_{s}(N^{\prime},P)\setminus\Phi_{s}(N,P)\,\Big{]}\,\Longrightarrow\,\big{[}\,i\in\Phi_{s}(N^{\prime},P)\,\big{]}.\quad$
Weak WrARP allows us to construct schools’ choice functions which are consistent with a priority order (Chambers and Yenmez, 2018). Intuitively, the property considers a situation where two students, $i$ and $j$, can be chosen from two different sets. If $i$ is chosen but not $j$ when both are available, then whenever $j$ is chosen, $i$ should also be chosen if she is available.
@@
Given a priority profile $\succ$, let ${\rm{DA}}^{\succ}$ be the mechanism that associates to each $(N,P)\in{\mathcal{N}}\times\mathcal{P}$ the student-optimal stable matching of $[N,S,\succ^{N},q,(P_{i})_{i\in N}]$, where $\succ^{N}$ is the priority profile induced by $\succ$. [^13] That is, if $\succ=(\succ_{s})_{s\in S}$, the induced priority profile $\succ^{N}=(\succ^{N}_{s})_{s\in S}$ is such that, for all $i,j\in N$ and $s\in S$, $i\succ^{N}_{s}j$ whenever $i\succ_{s}j$. When each school has only one available seat, Ehlers and Klaus (2016) demonstrate that a mechanism satisfies individual rationality, weak non-wastefulness, population-monotonicity, and strategy-proofness if and if it is equal to ${\rm{DA}}^{\succ}$ for some priority profile $\succ$. Furthermore, they show that this characterization does not hold in the general case in which schools have non-unit capacities.
The following result shows that the four axioms of Ehlers and Klaus (2016) jointly with weak WrARP and weak local non-bossiness fully characterize the family of mechanisms ${\rm{DA}}^{\succ}$, where $\succ$ is an arbitrary priority profile. Notice that the two new axioms are trivially satisfied when each school has only one available seat.
@@
A mechanism $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ is individually rational, weakly non-wasteful, population-monotonic, strategy-proof, weakly WrARP, and weakly locally non-bossy if and only if $\Phi={\rm{DA}}^{\succ}$ for some priority profile $\succ$.
The proof of this result and examples that show that the six axioms are independent are provided in the Appendix B.
In our characterization of DA, weak local non-bossiness is crucial to manage the schools with more than one seat available. Indeed, Example 1 of Ehlers and Klaus (2016), which we reproduce below, allows us to describe a mechanism that satisfies all the axioms of Theorem 2 but weak local non-bossines.
@@
Let $\overline{N}=\{1,2,3,4\}$, $S=\{s_{1},s_{2}\}$, and $(q_{s_{1}},q_{s_{2}})=(2,1)$. Consider the priority orders $\succ=(\succ_{s_{1}},\succ_{s_{2}})$ and $\succ^{\prime}=(\succ^{\prime}_{s_{1}},\succ_{s_{2}})$ characterized by $\succ_{s_{1}}:1,2,3,4$, $\succ_{s_{2}}:1,2,3,4$, and $\succ^{\prime}_{s_{1}}:1,2,4,3$. Denote by $top(P_{i})$ the most preferred alternative of student $i$ when her preferences are $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{N}}\times{\mathcal{P}}\rightarrow{\mathcal{M}}^{*}$ be the mechanism such that:
$\Omega(N,P)=\left\{\begin{array}[]{ll}{\rm{DA}}^{\succ^{\prime}}(N,P),&\quad\quad\mbox{when}\,\,\{1,2\}\subseteq N\,\mbox{and}\,top(P_{1})=top(P_{2})=s_{2};\\ {\rm{DA}}^{\succ}(N,P),&\quad\quad\mbox{in another case}.\end{array}\right.$
Ehlers and Klaus (2016, Example 1) show that $\Omega$ is individually rational, weakly non-wasteful, population-monotonic, and strategy-proof. Furthermore, $\Omega$ is weakly WrARP, because for each profile of preferences the school $s_{1}$ chooses students from a set following a priority order.
However, $\Omega$ is weakly locally bossy. Indeed, when preferences are given by
@@
we have that $\Omega(\overline{N},P)={\rm{DA}}^{\succ}(\overline{N},P)=((1,s_{2}),(2,s_{1}),(3,s_{1}),(4,s_{0})).$ If $P^{\prime}_{2}:s_{2},s_{1},s_{0}$, $\Omega(\overline{N},(P^{\prime}_{2},P_{-2}))={\rm{DA}}^{\succ^{\prime}}(\overline{N},(P^{\prime}_{2},P_{-2}))=((1,s_{2}),(2,s_{1}),(3,s_{0}),(4,s_{1})).$ Therefore, the student $2$ can change her colleagues without modify her school. [^14] This argument also shows that $\Omega$ is not group strategy-proof. Indeed, the coalition $\{2,4\}$ can report preferences $(P^{\prime}_{2},P_{4})$ to improve the assignment of $4$ without hurting $2$. Hence, the same example allows us to show that group strategy-proofness is a critical axiom in Corollary 2. $\Box$
A priority profile $\succ$ has an Ergin-cycle when for some $s,s^{\prime}\in S$ and $i,j,k\in\overline{N}$ we have that $i\succ_{s}j\succ_{s}k\succ_{s^{\prime}}i$ and there are disjoint sets $N_{s},N_{s^{\prime}}\subseteq\overline{N}\setminus\{i,j,k\}$, with $|N_{s}|=q_{s}-1$ and $|N_{s^{\prime}}|=q_{s^{\prime}}-1$, such that $j$ has lower priority in $s$ than everyone in $N_{s}$ and $i$ has lower priority in $s^{\prime}$ than everyone in $N_{s^{\prime}}$. $\succ$ is acyclic when it does not have Ergin-cycles.
@@
A mechanism $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ is individually rational, weakly non-wasteful, population-monotonic, group strategy-proof, and weakly WrARP if and only if $\Phi={\rm{DA}}^{\succ}$ for some acyclic priority profile $\succ$.
@@
Group strategy-proofness ensures that strategy-proofness and non-bossiness hold (Papái, 2000). Hence, the Theorem 2 implies that any mechanism $\Phi$ satisfying the five axioms described above coincides with ${\rm{DA}}^{\succ}$, for some priority profile $\succ$. In particular, as ${\rm{DA}}^{\succ}$ is group strategy-proof, $\succ$ is acyclic (Ergin, 2002). Reciprocally, given an acyclic priority profile $\succ$, it follows from Ergin (2002, Theorem 1) that ${\rm{DA}}^{\succ}$ is group strategy-proof (and Theorem 2 ensures that ${\rm{DA}}^{\succ}$ satisfies the other axioms). ∎
@@ 5. Mechanisms under preferences over colleagues
Most of the literature in school choice assumes that students only care about the school to which they are assigned. However, it is unrealistic that students do not care about the assignment of others. In particular, it has been empirically documented that the identity of colleagues is an important consideration when students are choosing a school (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
When students have arbitrary preferences over the set of matchings, many of the results in the literature break down. Specifically, a stable matching may not exist, even when students only care about the identity of the other students assigned to her school (Echenique and Yenmez, 2007). One way to overcome this problem is to restrict the domain of preferences. For instance, assuming that students have school-lexicographic preferences, meaning that they are primarily concerned with the school to which they are assigned (Sasaki and Toda, 1996; Dutta and Massó, 1997). In this domain, some of the standard results are recovered, such as the existence of a stable matching. However, a stable and strategy proof mechanism may not exist (Duque and Torres-Martínez, 2023). One may wonder if there is a subdomain of school-lexicographic preferences where stability and strategy-proofness are compatible. In this section, we show that within the subdomain of school-lexicographic preferences where only the schoolmates matter, the local non-bossiness of ${\rm{DA}}$ implies that this rule induces a stable and strategy-proof mechanism.
@@
Given a school choice context $[N,S,\succ,q]$, a school-lexicographic preference for student $i\in N$ is a complete and transitive preference relation $\unrhd_{i}$ defined on ${\mathcal{M}}$ such that, for any pair of matchings $\mu,\eta\in{\mathcal{M}}$, the following properties hold:
@@
The first condition states that a student cannot be indifferent between two matchings where she is assigned to different schools. By the second condition, if a student prefers a matching $\mu$ over $\eta$, and she is assigned to different schools in these two matchings, then she should prefer every matching where she is assigned to the same school as in $\mu$ to every matching where she is assigned to the same school as in $\eta$.
Notice that, in the preference domain ${\mathcal{D}}$ no restrictions are imposed on the ranking of two matchings in which a student is assigned to the same school. Let $P(\unrhd)=(P_{i}(\unrhd))_{i\in N}\in{\mathcal{P}}$ be the preferences over schools induced by $\unrhd=(\unrhd_{i})_{i\in N}\in{\mathcal{D}}$ through the following rule: given $s,s^{\prime}\in S\cup\{s_{0}\}$, $s\,P_{i}(\unrhd)\,s^{\prime}$ if and only if there exist $\mu,\mu^{\prime}\in{\mathcal{M}}$ such that $\mu(i)=s$, $\mu^{\prime}(i)=s^{\prime}$, and $\mu\rhd_{i}\mu^{\prime}$. The second property of Definition 5 ensures that the preferences $P(\unrhd)$ are well-defined.
@@
School-lexicographic preferences have no effects on stability, because the school choice problems $[N,S,\succ,q,\unrhd]$ and $[N,S,\succ,q,P(\unrhd)]$ have the same set of stable matchings (Sasaki and Toda, 1996; Dutta and Massó, 1997; Fonseca-Mairena and Triossi, 2023). However, there are strong effects on incentives as there may not exist a stable and strategy-proof mechanism (Duque and Torres-Martínez, 2023).
@@
Given $[N,S,\succ,q]$, the domain of school-lexicographic preferences over colleagues $\mathcal{D}_{{\rm{c}}}\subseteq\mathcal{D}$ is the set of profiles $(\unrhd_{i})_{i\in N}$ such that $\mu\rhd_{i}\eta$ and $\mu(i)=\eta(i)=s$ imply that $\mu(s)\neq\eta(s)$.
@@
The next result shows that there is a stable and strategy-proof mechanism in $\mathcal{D}_{{\rm{c}}}$, which is the student optimal stable mechanism applied to $[N,S,\succ,q,P(\unrhd)]$.
@@
In any school choice context $(S,N,\succ,q)$, the mechanism $\overline{\rm{DA}}:\mathcal{D}_{{\rm{c}}}\rightarrow\mathcal{M}$ defined by $\overline{\rm{DA}}(\unrhd)={\rm{DA}}(P(\unrhd))$ is stable and strategy-proof.
@@
  • Case 1: $\overline{\rm{DA}}_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})\neq\overline{\rm{DA}}_{i}(\unrhd)$. The definitions of $\overline{\rm{DA}}$ and $P_{i}(\unrhd)$ ensure that
    ${\rm{DA}}_{i}(P_{i}(\unrhd^{\prime}),P_{-i}(\unrhd))\,\,P_{i}(\unrhd)\,\,{\rm{DA}}_{i}(P(\unrhd)).$
    This implies that ${\rm{DA}}$ is not strategy-proof in ${\mathcal{P}}$, contradiction.
@@
    ${\rm{DA}}_{i}(P_{i}(\unrhd^{\prime}),P_{-i}(\unrhd))$ $=$ ${\rm{DA}}_{i}(P(\unrhd))\,=\,s,$
    ${\rm{DA}}_{s}(P_{i}(\unrhd^{\prime}),P_{-i}(\unrhd))$$\neq$ ${\rm{DA}}_{s}(P(\unrhd)).$
    This implies that ${\rm{DA}}$ is locally bossy, a contradiction to Theorem 1.

### v5

@@ 3. Local non-bossiness of DA
This property implies that $\rm{DA}$} satisfies the converse of local non-bossiness: if the colleagues of some student do not change when her preferences are modified, then her school does not change either. $\Box$
Remark 2.
Throughout the paper, we implicitly assume that schools have responsive preferences (Roth and Sotomayor, 1990), in the sense that priority orders $(\succ_{s})_{s\in S}$ and capacities $(q_{s})_{s\in S}$ determine the relevant information about how schools choose students from a pool of applicants. This assumption is essential to ensure our main result. Indeed, in Appendix B we provide examples illustrating that the local non-bossiness of $\rm{DA}$} cannot be guaranteed when schools’ preferences are non-responsive, even in contexts in which $\rm{DA}$} is well-defined, stable, strategy-proof, and optimal for students (cf., Roth, 1984b; Hatfield and Milgrom, 2005). $\Box$
The $\rm{DA}$} mechanism is not the only one that is locally non-bossy and stable in $\mathcal{P}$}. Indeed, the school-optimal stable mechanism also satisfies these properties, because it is non-bossy (Afacan and Dur, 2017). Furthermore, the arguments made in the proof of Theorem 1 can be easily adapted to provide a direct proof that the school-optimal stable mechanism is locally non-bossy. Indeed, since this mechanism assigns each student to the worst school she can reach in a stable matching, it is sufficient to repeat the proof focusing on $\mu$-worsening cycles: tuples $(i_{1},\ldots,i_{r})$ of different students such that $\mu(i_{l})P_{i_{l}}\mu(i_{l+i})$ for all $l\in\{1,\ldots,r\}$ [modulo $r$].
Not all stable mechanisms are locally non-bossy. For instance, the school-median stable mechanism (Klaus and Klijn, 2006) does not satisfy this property (see 10).
It is well-known that $\rm{DA}$} is the only mechanism defined in $\mathcal{P}$} that is both stable and strategy-proof (Alcalde and Barberà, 1994). Since local group strategy-proofness is stronger than strategy-proofness, and local non-bossiness and strategy-proofness ensure local group strategy-proofness (Lemma 1), the following result is a direct consequence of Theorem 1.
@@ 4. A characterization of DA via local non-bossiness
In this section, we provide a characterization of DA using axioms that do not involve stability. Our result extends the characterization of DA given by Ehlers and Klaus (2016, Theorem 1) from one-to-one matching problems to many-to-one contexts.
To maintain Ehlers and Klaus’ (2016) context, we analyze a variable-population school choice model. Hence, given a population $\overline{N}$ of students and a set $S$ of schools with capacities $q=(q_{s})_{s\in S}$, we consider mechanisms that depend only on the set $N\subseteq\overline{N}$ of students in the market and their preferences $(P_{i})_{i\in N}$. To simplify notation, we consider ${\mathcal{P}}={\mathcal{L}}^{\overline{N}}$ as the preference domain, although only the preferences of the agents in the market will be relevant at the moment of determining a matching. Let $\mathcal{N}$} be the collection of non-empty subsets of $\overline{N}$.
Given $N\in{\mathcal{N}}$, let $\mathcal{M}(N)$ be the set of matchings compatible with $[N,S,q]$. That is, $\mathcal{M}(N)$ is the family of functions $\mu:N\rightarrow S\cup\{s_{0}\}$ such that $|\{i\in N:\mu(i)=s\}|\leq q_{s}$ for all $s\in S$. Let $\mathcal{M}^{*}=\bigcup_{N\in{\mathcal{N}}}\mathcal{M}(N)$. For each $i\in\overline{N}$ and $P_{i}\in{\mathcal{L}}$, let
$A(P_{i})=\{s\in S:sP_{i}s_{0}\}$
be the set of schools that are admissible under $P_{i}$. Moreover, denote by $P^{t}_{i}$ the strict linear order defined on $A(P_{i})$ such that, for every $s,s^{\prime}\in A(P_{i})$, it holds that $sP^{t}s^{\prime}$ if and only if $sP_{i}s^{\prime}$. In other words, $P^{t}_{i}$ is the truncation of $P_{i}$ at $s_{0}$.
A mechanism is a function $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ that associates to each $(N,P)\in{\mathcal{N}}\times\mathcal{P}$ a matching in ${\mathcal{M}}(N)$ that depends only on $N,S,q,(P^{t}_{i})_{i\in N}]$. In other words, the matching $\Phi(N,P)$ can only be affected by the set of schools and their capacities, by the set of students in the market, and by the way in which these students rank their admissible schools. This last assumption allows us to avoid making $\Phi(N,P)$ depend on the alternatives that students in $N$ consider inadmissible, or on the preferences of individuals in $\overline{N}\setminus N$.
@@
By definition, for every mechanism $\Phi:{\mathcal{N}}\times{\mathcal{P}}\rightarrow{\mathcal{M}}^{*}$ and $(N,P)\in{\mathcal{N}}\times{\mathcal{P}}$, the matching $\Phi(N,P)$ depends only on $N,S,q,(P^{t}_{i})_{i\in N}]$. Hence, given a preference profile $P\in{\mathcal{P}}$ such that $A(P_{i})=\{s\}$ for all $i\in{\overline{N}}$, the mapping that associates to each $N\subseteq\overline{N}$ the set $\Phi_{s}(N,P)$ can be viewed as a choice function for school $s$, in the sense that it determines the students assigned to $s$ for every subset of candidates in $\overline{N}$.
@@
A mechanism $\Phi$ satisfies the S-weaker axiom of revealed preference (S-WrARP) when for every $i,j\in\overline{N}$, $s\in S$, and $N,N^{\prime}\in{\mathcal{N}}$ such that $i,j\in N\cap N^{\prime}$ and $|N|=|N^{\prime}|=q_{s}+1$, the following property holds when $P\in{\mathcal{P}}$ satisfies $A(P_{k})=\{s\}$ for all $k\in\overline{N}$,
$\quad\Big[\,i\in\Phi_{s}(N,P),\,\,\quad j\in\Phi_{s}(N^{\prime},P)\setminus\Phi_{s}(N,P)\,\Big]\,\Longrightarrow\,\big[\,i\in\Phi_{s}(N^{\prime},P)\,\big].\quad$
S-WrARP is a restricted version of the weaker axiom of revealed preference introduced by Jamison and Lau (1973), but applied to schools’ choice functions induced by a mechanism. More precisely, assuming that some school $s$ is the only admissible alternative, consider a situation in which two students, $i$ and $j$, can be assigned to $s$ from two different sets of size $q_{s}+1$, namely $N$ and $N^{\prime}$. The property S-WrARP ensures that, when the mechanism $\Phi$ is implemented, if $i$, but not $j$, is assigned to $s$ when the set $N$ is considered, then whenever $j$ is assigned to $s$ when $N^{\prime}$ is considered, $i$ is also assigned to $s$. This axiom will allow us to appeal to the results of Chambers and Yenmez (2018) to construct schools’ priority orders that are consistent with the assignments determined by $\Phi$.
@@
Given a priority profile $\succ$, let $\rm{DA}$$^{\succ}$ be the mechanism that associates to each $(N,P)\in{\mathcal{N}}\times\mathcal{P}$ the student-optimal stable matching of $N,S,\succ^{N},q,(P_{i})_{i\in N}]$, where $\succ^{N}$ is the priority profile induced by $\succ$. [^15] That is, if $\succ=(\succ_{s})_{s\in S}$, the induced priority profile $\succ^{N}=(\succ^{N}_{s})_{s\in S}$ is such that, for all $i,j\in N$ and $s\in S$, $i\succ^{N}_{s}j$ whenever $i\succ_{s}j$. When each school has only one available seat, Ehlers and Klaus (2016) demonstrate that a mechanism $\Phi$ satisfies individual rationality, weak non-wastefulness, population-monotonicity, and strategy-proofness if and only if $\Phi={\rm{DA}}^{\succ}$ for some priority profile $\succ$. Furthermore, they show that this characterization does not hold in the general case in which schools have non-unit capacities.
The following result shows that the four axioms of Ehlers and Klaus (2016) jointly with S-WrARP and weak local non-bossiness fully characterize the family of mechanisms $\rm{DA}$$^{\succ}$, where $\succ$ is an arbitrary priority profile. Note that the two new axioms are trivially satisfied when each school has only one available seat.
@@
A mechanism $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ is individually rational, weakly non-wasteful, population-monotonic, strategy-proof, weakly locally non-bossy, and satisfies S-WrARP if and only if $\Phi={\rm{DA}}^{\succ}$ for some priority profile $\succ$.
The proof of this result and examples that show that the six axioms are independent are provided in Appendix C.
In our characterization of DA, weak local non-bossiness is crucial to manage the schools with more than one seat available. Indeed, Example 1 of Ehlers and Klaus (2016)—which we reproduce below—allows us to describe a mechanism that satisfies all the axioms of Theorem 2 but weak local non-bossiness.
@@
Let $\overline{N}=\{1,2,3,4\}$, $S=\{s_{1},s_{2}\}$, and $(q_{s_{1}},q_{s_{2}})=(2,1)$. Consider the priority orders $\succ=(\succ_{s_{1}},\succ_{s_{2}})$ and $\succ^{\prime}=(\succ^{\prime}_{s_{1}},\succ_{s_{2}})$ characterized by $\succ_{s_{1}}:1,2,3,4$, $\succ_{s_{2}}:1,2,3,4$, and $\succ^{\prime}_{s_{1}}:1,2,4,3$. Denote by $top(P_{i})$ the most preferred alternative of a student $i$ when her preferences are $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{N}}\times{\mathcal{P}}\rightarrow{\mathcal{M}}^{*}$ be the mechanism such that:
$\Omega(N,P)=\left\{\begin{array}[]{ll}{\rm{DA}}^{\succ^{\prime}}(N,P),&\quad\quad\mbox{when}\,\,\{1,2\}\subseteq N\,\mbox{and}\,top(P_{1})=top(P_{2})=s_{2};\\ {\rm{DA}}^{\succ}(N,P),&\quad\quad\mbox{otherwise}.\end{array}\right$.
Ehlers and Klaus (2016, Example 1) show that $\Omega$ is individually rational, weakly non-wasteful, population-monotonic, and strategy-proof. Furthermore, $\Omega$ satisfies S-WrARP, because for each profile of preferences schools choose students from a set following a priority order. However, $\Omega$ is weakly locally bossy. Indeed, when preferences are given by
@@
we have that $\Omega(\overline{N},P)={\rm{DA}}^{\succ}(\overline{N},P)=((1,s_{2}),(2,s_{1}),(3,s_{1}),(4,s_{0})).$ If $P^{\prime}_{2}:s_{2},s_{1},s_{0}$, $\Omega(\overline{N},(P^{\prime}_{2},P_{-2}))={\rm{DA}}^{\succ^{\prime}}(\overline{N},(P^{\prime}_{2},P_{-2}))=((1,s_{2}),(2,s_{1}),(3,s_{0}),(4,s_{1})).$ Therefore, the student $2$ can change her colleagues without modifying her school. [^16] This argument also shows that $\Omega$ is not group strategy-proof. Indeed, the coalition $\{2,4\}$ can report preferences $(P^{\prime}_{2},P_{4})$ to improve the assignment of $4$ without hurting $2$. Hence, the same example allows us to show that group strategy-proofness is a critical axiom in Corollary 2. $\Box$
A priority profile $\succ$ has an Ergin-cycle when for some $s,s^{\prime}\in S$ and $i,j,k\in\overline{N}$ we have that $i\succ_{s}j\succ_{s}k\succ_{s^{\prime}}i$ and there are disjoint sets $N_{s},N_{s^{\prime}}\subseteq\overline{N}\setminus\{i,j,k\}$, with $|N_{s}|=q_{s}-1$ and $|N_{s^{\prime}}|=q_{s^{\prime}}-1$, such that $j$ has lower priority in $s$ than everyone in $N_{s}$ and $i$ has lower priority in $s^{\prime}$ than everyone in $N_{s^{\prime}}$. The priority profile $\succ$ is acyclic when it does not have Ergin-cycles.
@@
A mechanism $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow\mathcal{M}^{*}$ is individually rational, weakly non-wasteful, population-monotonic, group strategy-proof, and satisfies S-WrARP if and only if $\Phi={\rm{DA}}^{\succ}$ for some acyclic priority profile $\succ$.
@@
Group strategy-proofness ensures that strategy-proofness and non-bossiness hold (Papái, 2000). Hence, Theorem 2 implies that a mechanism $\Phi$ satisfying the five axioms described above coincides with $\rm{DA}$$^{\succ}$, for some priority profile $\succ$. In particular, as $\rm{DA}$$^{\succ}$ is group strategy-proof, $\succ$ is acyclic (Ergin, 2002). Reciprocally, given an acyclic priority profile $\succ$, it follows from Ergin (2002, Theorem 1) that $\rm{DA}$$^{\succ}$ is group strategy-proof (and Theorem 2 ensures that $\rm{DA}$$^{\succ}$ satisfies the other axioms). ∎
@@ 5. Mechanisms under preferences over colleagues
Most of the literature on school choice assumes that students only care about the school to which they are assigned. However, it is unrealistic that students do not care about the assignment of others. In particular, it has been empirically documented that the identity of colleagues is an important consideration when students are choosing a school (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
When students have arbitrary preferences over the set of matchings, many of the results in the literature break down. Specifically, a stable matching may not exist, even when students only care about the identity of the other students assigned to her school (Echenique and Yenmez, 2007). One way to overcome this problem is to restrict the domain of preferences. For instance, suppose that students have school-lexicographic preferences, meaning that they are primarily concerned with the school to which they are assigned (Sasaki and Toda, 1996; Dutta and Massó, 1997). In this domain, some of the standard results are recovered, such as the existence of a stable matching. However, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). One may wonder if there is a subdomain of school-lexicographic preferences where stability and strategy-proofness are compatible. In this section, we show that within the subdomain of school-lexicographic preferences where only the schoolmates matter, the local non-bossiness of $\rm{DA}$} implies that this rule induces a stable and strategy-proof mechanism.
@@
Given a school choice context $N,S,\succ,q]$, a school-lexicographic preference for student $i\in N$ is a complete and transitive preference relation $\unrhd_{i}$ defined on $\mathcal{M}$} such that, for every pair of matchings $\mu,\eta\in{\mathcal{M}}$, the following properties hold:
@@
The first condition states that a student cannot be indifferent between two matchings where she is assigned to different schools. By the second condition, if a student prefers a matching $\mu$ over $\eta$, and she is assigned to different schools in these matchings, then she should prefer every matching where she is assigned to the same school as in $\mu$ to every matching where she is assigned to the same school as in $\eta$.
Notice that, in the preference domain $\mathcal{D}$} no restrictions are imposed on the ranking of two matchings in which a student is assigned to the same school. Let $P(\unrhd)=(P_{i}(\unrhd_{i}))_{i\in N}\in{\mathcal{P}}$ be the preferences over schools induced by $\unrhd=(\unrhd_{i})_{i\in N}\in{\mathcal{D}}$ through the following rule: given $s,s^{\prime}\in S\cup\{s_{0}\}$, $s\,P_{i}(\unrhd_{i})\,s^{\prime}$ if and only if there exist $\mu,\mu^{\prime}\in{\mathcal{M}}$ such that $\mu(i)=s$, $\mu^{\prime}(i)=s^{\prime}$, and $\mu\rhd_{i}\mu^{\prime}$. The second property in Definition 5 ensures that $P(\unrhd)$ is well-defined.
@@
School-lexicographic preferences have no effect on stability, because the school choice problems $N,S,\succ,q,\unrhd]$ and $N,S,\succ,q,P(\unrhd)]$ have the same set of stable matchings (Sasaki and Toda, 1996; Dutta and Massó, 1997; Fonseca-Mairena and Triossi, 2023). However, there are strong effects on incentives as there may not exist a stable and strategy-proof mechanism (Duque and Torres-Martínez, 2023).
@@
Given $N,S,\succ,q]$, the domain of school-lexicographic preferences over colleagues $\mathcal{D}_{{\rm{c}}}\subseteq\mathcal{D}$ is the set of profiles $(\unrhd_{i})_{i\in N}$ such that $\mu\rhd_{i}\eta$ and $\mu(i)=\eta(i)=s$ imply that $\mu(s)\neq\eta(s)$.
@@
The next result shows that there is a stable and strategy-proof mechanism in $\mathcal{D}_{{\rm{c}}}$, which is the student-optimal stable mechanism applied to $N,S,\succ,q,P(\unrhd)]$.
@@
In every school choice context $(S,N,\succ,q)$, the mechanism $\overline{\rm{DA}}:\mathcal{D}_{{\rm{c}}}\rightarrow\mathcal{M}$ defined by $\overline{\rm{DA}}(\unrhd)={\rm{DA}}(P(\unrhd))$ is stable and strategy-proof.
@@
  • Case 1: $\overline{\rm{DA}}_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})\neq\overline{\rm{DA}}_{i}(\unrhd)$. The definitions of $\overline{\rm{DA}}$ and $P_{i}(\unrhd_{i})$ ensure that
    ${\rm{DA}}_{i}(P_{i}(\unrhd^{\prime}_{i}),P_{-i}(\unrhd))\,\,P_{i}(\unrhd_{i})\,\,{\rm{DA}}_{i}(P(\unrhd)).$
    This implies that $\rm{DA}$} is not strategy-proof in $\mathcal{P}$}. A contradiction.
@@
    ${\rm{DA}}_{i}(P_{i}(\unrhd^{\prime}_{i}),P_{-i}(\unrhd))$ $=$ ${\rm{DA}}_{i}(P(\unrhd))\,=\,s,$
    ${\rm{DA}}_{s}(P_{i}(\unrhd^{\prime}_{i}),P_{-i}(\unrhd))$ $\neq$ ${\rm{DA}}_{s}(P(\unrhd)).$
    This implies that $\rm{DA}$} is locally bossy, which contradicts Theorem 1.